durin the first part of a trip a canoesit travels 64 miles at a certain speed. The conoeist treavels 22 miles on the secondo part of the trip at a speed of 5 mph slower. the total time for the trip is 4 hours what was the speed of each part of the trip.
Answer by mananth(16946)

durin the first part of a trip a canoesit travels 64 miles at a certain speed. The conoeist treavels 22 miles on the secondo part of the trip at a speed of 5 mph slower. the total time for the trip is 4 hours what was the speed of each part of the trip.
..
let speed be x mph during the first part
distance = 64 miles
time taken = 64/x
...
second part
speed = x-5
distance = 22 miles.
time = 22/(x-5)
..
time first leg + time second leg = total time
64/x+22/(x-5)=4
64(x-5)+22x/ x(x-5)=4
64x-320+22x= 4x(x-5)
86x-320=4x^2-20x
4x^2-86x-20x+320=0
4x^2-106x+320=0
/2
2x^2-53x+160=0
quadratic equation to find the roots.
discriminant = 1529
x1= (53+sqrt(1529))/4
x1=23.02
x2=(53-sqrt(1529))/4
x2= 3.47 ( not possible)
first leg speed = 23mph
second leg = 18 mph
...
m.ananth@hotmail.ca